OpenStudy (anonymous):

2logx – log4 = 0 need help with this

OpenStudy (anonymous):

Rearrange:\[\bf 2\log(x)=\log(4)\]Make the 2 the exponent of x:\[\bf \log(x^2)=\log(4)\]Both sides have a single logarithm with the same base which implies that:\[\bf \implies x^2 = 4 \implies x = \pm 2\]But -2 would be an extraneous solution since you cannot have the logarithm of a negative number so we eliminate that solution and we are left with:\[\bf x = 2\] @raevenbennett